OpenStudy (anonymous):

Let f(x) = x + 6 and g(x) = x - 5. What is f(x) • g(x)? A) x2 + 6x + 30 B) x2 - 6x + 1 C) x2 - x + 30 D) x2 + x - 30

OpenStudy (redprince):

f(x).g(x)=x^2+x-30

OpenStudy (anonymous):

f(x) = x^2 + 3x − 10 and g(x) = x − 2 (i) f(x) + g(x) =x^2+3x-10+x-2 =x^2+4x-12 answer// (ii) 2g(x) − f(x) ..=2(x-2)-(x^2+3x-10) =2x-4-x^2-3x+10 =-x^2-x+6 answer// (iii)g(x)/f(x) ..x-2 =------------ x^2+3x-10 ....x-2 =----------- ..(x+5)(x-2) .....1 =--------- answer// ...x+5 source: http://nz.answers.yahoo.com/question/index?qid=20100324023239AAlAlgo

OpenStudy (aravindg):

hehe

OpenStudy (aravindg):

yahoo answers!

OpenStudy (redprince):

D is the right answer I solve it myself!

OpenStudy (redprince):

ty